### v1

# Directional and contra-directional coupling in Huygens’ metawaveguide microring resonators

Directional and contra-directional coupling in Huygens’ metawaveguide microring resonators
††journal: opticajournal††articletype: Research ArticleHuygens’ metawaveguides represent a transformative concept in photonic device engineering, enabling unprecedented control over light propagation. This study presents, for the first time, integrated Huygens’-based microring resonators and directional and contra-directional couplers, specifically designed for operation at the 1550 nm telecommunication wavelength. By leveraging the unique properties of resonant Huygens’ waveguides, we demonstrate efficient evanescent directional coupling with high- resonators, characterized by negative group index and near-zero dispersion, which are critical for enhancing performance in compact, high-performance add-drop filters. The research further explores the implications of these novel structures on group index and group velocity dispersion, providing insights into their potential applications in nonlinear optics and quantum information technologies. Notably, the introduction of a hybrid subwavelength grating-Huygens’ contra-directional coupler facilitates backward coupling between resonant and non-resonant metawaveguides, achieving a broad spectral rejection bandwidth. Our findings advance the integration of resonant metamaterials into scalable photonic platforms, laying the groundwork for innovative applications in optical communications, quantum photonics and sensing systems.
1 Introduction
Photonic metamaterials have fundamentally transformed our capacity to manipulate both free-space and guided optical waves, enabling functionalities far beyond those achievable with conventional dielectric media. Through precise engineering of their optical parameters, metamaterials provide unprecedented control over key properties of light—such as amplitude, phase, polarization, wavelength, and group velocity [1, 2, 3, 4, 5, 6, 7]. This control is typically realized by subwavelength-scale periodic modulation of materials with contrasting refractive indices [7, 8, 9, 10, 11, 12, 13]. Moreover, by harnessing electromagnetic resonances within these periodic nanostructures, the light–matter interaction can be significantly amplified [14, 15, 16, 17].
Resonant metamaterials exhibit a range of distinctive functionalities, including tailored light transmission [18], enhanced absorption [19], control over propagation direction [20, 21], beam shaping, focusing and lensing [22, 23], sensing [24], and enhancement of nonlinear optical effects [25]. These capabilities are vital for diverse applications spanning classical and quantum information processing, communication, and signal routing [26, 27, 28].
Manipulation of the effective refractive index and group index is essential for optimizing mode coupling [29, 30] and managing pulse propagation and dispersion in integrated photonics [31, 32, 33]. Such control enables advanced signal processing techniques indispensable to high-performance optical networks. Dispersion-engineered metamaterials, especially those incorporating subwavelength grating (SWG) structures that modulate the effective index [34, 35, 36], have emerged as key components in fiber-optic communication systems. These structures facilitate efficient fiber-to-chip coupling [37, 38],mode conversion [39], wavelength multiplexing [37], and high-performance directional and contra-directional couplers [40, 41, 42, 43, 44, 45], as well as micro-resonators [46, 47, 48, 49] and other advanced photonic devices [12, 13].
In parallel, recent studies have unveiled remarkable optical phenomena — such as backward pulse propagation [50, 51, 52, 53, 54, 55], pulse compression [56, 57, 58, 59], ultrafast signal processing and switching [60, 61, 62], and soliton and frequency-comb generation [63, 64, 65] — enabled by waveguides exhibiting negative group index and anomalous dispersion [66, 67, 68]. Negative group index, wherein the phase velocity is opposite to the energy flow, gives rise to counterintuitive phenomena such as reversed pulse propagation and enhanced control of optical information. These regimes are accessible in resonant metamaterials [69, 70].
While early implementations of resonant metawaveguides based on periodic Mie-scattering nanoparticles were hindered by excessive scattering and absorption losses [7, 70, 14, 71], recent progress has led to the realization of low-loss Huygens’ metawaveguides [72, 73]. These structures exploit Huygens’ particles, which achieve backscattering suppression through constructive interference between co-aligned electric dipole (ED) and magnetic dipole (MD) resonances—satisfying the Kerker condition [74, 75, 18, 76, 77]. This condition ensures forward-directed radiation, enabling light propagation via near-field coupling between adjacent Huygens’ particles, while preventing scattering into radiation modes.
In this work, we present Huygens’ metawaveguide directional couplers and ring resonators designed to operate at the 1550 nm telecommunication wavelength. These structures are used to characterize the group index and group velocity dispersion of resonant metawaveguides and to assess the performance of Huygens’ metawaveguide ring resonators as compact, high-performance add-drop filters. We further demonstrate efficient evanescent directional coupling between Huygens’ waveguides and high- resonators engineered from metawaveguides with negative group index, near-zero dispersion, and anomalous dispersion profiles. Finally, we introduce a novel SWG–Huygens’ contra-directional coupler, enabling backward coupling between resonant and non-resonant metawaveguides while achieving a broad spectral rejection bandwidth.
2 Results and Discussions
2.1 Huygens’ nanoantennas and metawaveguides
The core components of Huygens’ metawaveguides are individual dielectric nanoantennas that simultaneously exhibit Mie resonant electric and magnetic dipole (ED and MD) scattering within a desired spectral range [72]. In Fig. 1a(i) the ED and MD resonances in a Huygens particle are shown schematically, with the respective moments and satisfying the forward Kerker condition, (where is the medium’s relative permittivity and is the speed of light in vacuum), ensuring their comparable magnitudes and spectral overlap [78, 79]. This in-phase excitation creates constructive interference of the ED and MD scattering in the forward direction and destructive interference in the backward direction, hence in zero backward scattering. Under this condition, the nanoantennas function as forward directional Huygens’ scatterers.
Through multipolar decomposition analysis of the total scattering cross-section [80] (see Fig. S1 in Supplement 1), we identified that single-crystal silicon nano-cuboid antennas with dimensions nm, nm, and nm embedded in a silica cladding act as unidirectional forward Huygens’ scatterers at the 1550 nm telecommunication wavelength. This is confirmed by the near-field mode profile shown in Fig. 1a(ii). The height nm was specifically chosen to ensure compatibility with the standard CMOS compatible silicon-on-insulator (SOI) waveguide platform. We carefully tuned the cuboid width-to-length ratio, , for TE polarization to achieve spectral overlap of the ED and MD resonances near the 1550 nm wavelength. We carried out the multipolar expansion of the individual Huygens’ nanoantennas by exciting it with a totalfield scattered-field (TFSF) source in a commercial software Ansys Lumerical FDTD, and performed the decomposition of the total scattering cross-section using the generalized Kerker condition calculation.
Next, we employed the optimized Huygens’ antennas to construct a Huygens’ metawaveguide, setting the antenna periodicity, , and the antenna length, , and introducing an inter-antenna gap, (see Fig. 1(b)). The gap was optimized within the range 80-140 nm to maximize the combined forward-directional scattering efficiency of a dimer, composed of two consecutive nanoantennas. In this range, the fields of adjacent antennas remain sufficiently decoupled to preserve the fundamental scattering properties of individual antennas, while the proximity ensures the second antenna effectively captures the light scattered from the first antenna, enabling waveguiding effect.
The calculated dispersion band of the infinitely long metawaveguide (see Fig. 1(c)) reveals two guided modes separated by the photonic band gap (PBG) around 1600 nm. The light line of SiO2 is indicated in yellow. The fundamental guided TE mode lies primarily in the long-wavelength region (below the PBG), referred to as the SWG band. Here, the chosen nanoantenna dimensions are considerably smaller than the wavelength of the propagating light. The SWG band extends up to the PBG at the edge of the Brillouin zone, where the array period equals half the effective wavelength. Below the PBG, the behavior of the SWG mode can be described by expressing the effective material refractive index, , as a weighted average of the constituent refractive indices. The second guided mode, positioned above the PBG near the Huygens’ nanoantenna resonance wavelength, operates in the Huygens’ band and exhibits a negative group index [72].
We chose the antenna periodicity, = 430 nm, to spectrally position the Huygens’ band in the 1550 nm telecommunication wavelength range and calculated the Bloch modes for a discrete set of wavevectors in normalized ranging from 0.3 to 0.5 (2). The chosen periodicity, of individual antennas with width, constitutes the duty cycle, = 0.73. Next, we calculated the field profiles of the Bloch modes ( from the middle of the – cross-section of the antenna) of each branch at = 0.425 and 1550 nm for the chosen periodicity, and launch the Bloch mode through the straight Huygens’ metawaveguide. We performed a multipolar decomposition analysis of the fields excited inside the individual antennas for the 1D metawaveguide periodic array. Multipole decomposition for the Huygens’ band revealed a good spectral overlap between ED and MD resonances, as required to fulfill the Kerker condition over a broadband (see Fig. S1 in Supplement 1). The resonant properties of the coupled individual antennas are preserved in the array while the relative magnitudes of the electric and magnetic moments are locked across the Huygens band. The antennas in the array effectively act as coupled Huygens’ scatterers over a wide spectral range of 60 nm [72].
Following the same excitation technique described above, we performed the simulations of the Huygens’ directional couplers and the micro-resonators using the Ansys Lumerical FDTD software. In addition, we performed band diagram simulations of the SWG–Huygens’ contra-directional couplers using the MIT Photonic Bands (MPB) solver. Next, we fabricated the samples using a commercial foundry service, Applied Nanotools Inc. Fabrication involved electron-beam lithography on a commercial SOI substrate, comprising a 220 nm-thick single-crystalline silicon layer above a 3m buried oxide (BOX) layer on a silicon handle wafer. Electron-beam lithography was used to define waveguide patterns, which were subsequently transferred into the silicon layer through plasma etching. The samples were then coated with a 3 m SiO2 cladding layer using plasma-enhanced chemical vapor deposition (PECVD). A second lithographic step and etching were performed to create the optical facets for light coupling, involving vertical etching through the top oxide, BOX layers, and into the silicon substrate. Lastly, the chips were diced near the etched facets.
The fabricated samples were characterized using a continuous wave (CW) laser Agilent 8164B Lightwave Measurement System (LMS) with a linewidth of 0.01 nm operating in the wavelength range of 1460 nm to 1600 nm. To launch the input and collect and output light to measure the transmission, we used two spherical lensed fibers (radius of curvature: = 9 m and mode field diameter: MFD = 3 m); one was connected to the TE polarization-maintained laser source output, and another to the power meter. Every time we measured the transmission, we scanned the CW laser across its operating wavelength range to cover the transmission spectra of the ring resonators. The input and outputs of the bus waveguides were directly connected to the on-chip fiber-to-edge couplers [37] to couple the source light and collect the transmitted signal.
2.2 Directional coupling of Huygens’ metawaveguides
We designed and fabricated on-chip SOI directional couplers by placing two straight Huygens’ metawaveguides in parallel, separated by a coupling gap, , to enable evanescent mode coupling, and finally measured the transmission power over the entire Huygens’ band from 1480 nm to 1560 nm wavelength for all the devices. The mode propagation in the directional coupler and the corresponding scanning electron microscopy (SEM) image of the device are shown in Fig.1(d) and Fig.1(e), respectively. The coupling length was designed as m and the duty cycle to to achieve 100 coupling efficiency [46].
To evaluate the impact of the coupling gap on the propagation characteristics of the coupled Huygens’ metawaveguides, we conducted a multipolar expansion analysis of individual antennas at the central location of the coupler. Fig.1(f) presents the simulated scattering cross-section for decomposed multipolar components with a minimum coupling gap of 100 nm. While the electric dipole (ED) and magnetic dipole (MD) are excited in phase near 1576 nm, they exhibit an amplitude mismatch of nearly a factor of 2, yet remain within the same order of magnitude. The amplitude matching is improved for increased coupling gaps as shown in Fig.1(g), while the spectral overlap remains consistent (see Fig. S1 in Supplement 1). Fig. 1(h) reveals that the coupling efficiency of the straight Huygens’ directional couplers strongly depends on the gap between the adjacent waveguides. Experimental results arepresented in Fig. S2 in Supplement 1. This behavior is expected since variations in the gap affect significantly the mode index of the Huygens’ waveguide in the coupling region.
Since Huygens’ scatterers are designed to suppress back-scattering (see Fig.1(a)), they effectively reduce propagation losses associated with back-reflections in periodic resonant metawaveguides. Our results confirm that back-reflection remains suppressed even in coupled Huygens’ metawaveguides, where energy exchange between grating structures could otherwise induce back-scattering [81]. Comparisons with a non-resonant SWG metawaveguide directional coupler with the same coupling length ( = 9 m), duty cycle ( = 0.72), antenna width ( = 180 nm), and periodicity ( = 250 nm) show that backward directional coupling remains lower in the coupled Huygens’ waveguides compared to the SWG waveguides, as shown in Fig.1(i).
2.3 Huygens’ ring resonators
Next, we substituted one of the two straight Huygens’ metawaveguides of the directional coupler with a curved metawaveguide. To investigate the cross-coupling characteristics, we analyzed such straight-bend metawaveguide directional couplers by varying the coupling gap, while keeping the bend radius fixed at m to minimize bend loss. Fig.2(a) depicts the simulated mode propagation of the directional coupler with a coupling gap of 150 nm, and Fig.2(b) shows a corresponding SEM image of a fabricated structure. As shown in Fig. 2(c), unlike the straight directional couplers, the cross-coupling efficiency decreases monotonously as the gap increases. This is expected since coupled light power cannot be efficiently exchanged between the straight and curved metawaveguides without a sufficiently long coupling section.
We utilized the straight-bend metawaveguide directional coupler to construct a Huygens’ ring resonator. Fig.2(d) shows an SEM image of the fabricated ring resonator with the radius of 9 m coupled to a bus metawaveguide via a gap of 200 nm. The inset shows the calculated resonator electric field at a 1550 nm wavelength. The transmission spectra near the Huygens’ band, presented in Fig.2(e), reveal that the free-spectral range (FSR), , of the excited resonance modes remains almost constant in the central spectral zone, but gradually decreases as the wavelength approaches the photonic bandgap (PBG). The measured -factor of the resonances across the Huygens’ band is around 2000.
Using the FSR corresponding to the resonances shown in Fig.2(e), we extracted the metawaveguide group index, , from the relation , where is the speed of light in vacuum and is the length of the ring. Due to the inverse frequency-wavevector dispersion relation (shown in Fig.1(c)), the group index, , for the Huygens’ band is negative. The flat and descending FSR depicted in Fig.2(e) correspond to zero and anomalous group velocity dispersion (GVD), respectively. The dispersion parameter, , where , also follows this behavior. Fig.2(f) depicts the group index, , averaging -4.65 in the central zone of the Huygens’ band, decreasing to -10 near the PBG edges. In contrast, the dispersion parameter, , in the same plot demonstrates near zero dispersion in the central zone which eventually declines to -10,000 ps/nm·km as the photonic bandgap is approached. Such feature is unique for the Huygens’ metawaveguide ring resonators, as shown in the comparison with the conventional and the SWG ring resonators presented in Fig. S3 in Supplement 1.
2.4 Characteristics of Huygens’ micro-resonator add-drop filters
We also studied four-port add-drop filter resonators using the Huygens’ directional couplers demonstrated earlier. Figures 3(a) and 3(b) present SEM images of the fabricated ring and racetrack resonators, each featuring input, through, add, and drop ports. In both designs, the radii of the curved sections are set to = 9 m, while the straight coupling length of the racetrack resonators is = 9 m. Figures 3(b-d) show the experimentally observed transmission spectra at the through, drop, and add ports near the Huygens’ band for coupling gaps of 150 nm, 200 nm, and 250 nm in the ring resonator filters. Similarly, Figures 3(e-g) present the corresponding spectra for the racetrack resonators.
In the ring-based filters, the observed resonances exhibit over-coupled, critically coupled, and under-coupled characteristics as the coupling gap increases. The extinction ratio of the through-port resonances gradually decreases with increasing coupling gap, as the coupling strength is reduced. For both over- and critically coupled regimes, the insertion loss difference between the add and drop ports is approximately 10 dB. In the under-coupled regime, some resonances exhibit spectral splitting [82, 81, 83], accompanied by a reduced insertion loss difference between the add and drop ports. The resonance splitting would become more pronounced for larger coupling gaps, as shown in Fig. S4 in Supplement 1.
In the racetrack filter, the resonances transition from over-coupled to critically coupled and then to under-coupled as the coupling gap increases. For the smallest coupling gap, the over-coupled resonance extinction ratio at the through port is approximately 5–10 dB. This value increases to 30–40 dB for critically coupled resonances when the gap is increased by about 50 nm. With an additional 50 nm increment, under-coupling becomes pronounced and the through-port transmission drops to 20 dB below the drop-port transmission. Unlike the ring resonators, here the resonances are split in the over-coupled regime with the spectral separation exceeding the 3-dB bandwidth. Consequently, the insertion loss at the add and drop ports becomes comparable. The asymmetric split resonances suggest back-coupling effects in the Huygens’ racetrack resonators [82, 81], which may arise from the waveguide bends or the directional coupler sections.
Figure 3(i) illustrates the measured -factors of ring resonances at the 3-dB bandwidth, which progressively increase with the coupling gap. The average -factor remains around 103. Conversely, Figure 3(j) reveals an opposing trend for racetrack split resonances, where the -factors for the smallest coupling gaps range between 104 and 105—approximately one to two orders of magnitude higher than those observed for larger gaps. This significant enhancement in the -factor arises from the low coupling ratio, which means that the -factor approaches the intrinsic of an unperturbed ring.
2.5 Add-drop filtering with contra-directional couplers
Contra-directional couplers (CDC) are optical metamaterial structures that enable wavelength-specific power exchange between different waveguide supermodes traveling in opposite directions. They function as wavelength filters, with the backward-coupled fields directed to a port distinct from the input. When integrated into silicon photonics platforms, CDCs provide a compact and effective solution for on-chip spectral control of light. In this section of the article, we design and experimentally demonstrate a CDC by side-loading the Huygens’ waveguide with a modulated subwavelength grating (SWG) waveguide, as shown in Fig. 4(a). The use of a SWG waveguide enables considerable interaction between the supermodes of the structure while preserving the radiation pattern of the Huygens’ waveguide segments. We then experimentally assess optical add-drop filters comprising both a Huygens’ directional coupler and a Huygens’ CDC, demonstrating that the latter provides spectral control over the racetrack resonances.
To design the CDC, we first set the period of the SWG waveguide to = 215 nm, which is precisely half the period of the Huygens’ waveguide ( = 430 nm), as shown in Fig. 4(b). The duty cycle of the SWG waveguide was set to 0.5, ensuring equal dimensions for both the silicon blocks and gaps ( 108 nm). This balance facilitates fabrication by avoiding overly small features. The optimal center-to-center separation between the Huygens’ waveguide and the SWG waveguide was determined to be = 850 nm, maximizing evanescent field coupling while conserving the radiation pattern of the Huygens’ waveguide segments.
By initially setting the width of the SWG waveguide segments equal to the width of the Huygens’ waveguide segments, we observed a small bandgap of approximately 1.5 nm at the center of the Huygens’ waveguide operating bandwidth ( = 1550 nm). This bandgap corresponds to the back-coupling between the Huygens’ waveguide and the SWG waveguide supermodes and was analyzed using band diagram calculations, as exemplified in Fig. 4(c). To increase the bandgap width, we narrowed one segment in each pair of silicon segments in the SWG waveguide, we made the inner segment width () smaller. This change induced a blue-shift in the bandgap center wavelength, attributed to the reduced silicon content thus a lower equivalent refractive index in the SWG waveguide. To shift the bandgap center back to the target wavelength ( = 1550 nm), we optimally widened the adjacent segments, we made the outer segment width ( ) larger. After repeating this procedure for each inner segment width in the range [0.2,0.55] m, we obtained the simulation results shown in Fig. 4(d). The maximum achievable bandgap was found to be of approximately 15 nm. Both the bandgap as a function of inner segment width, (), and the outer segment width as a function of inner segment width, (), were fitted with second-degree polynomials, providing the basis for the design of an apodized filter.
Building on the previous bandgap calculations, we designed an apodized single-band filter with a peak bandwidth of = 10 nm. The length of the CDC was set to = 150 m, and a Gaussian apodization window was applied, with the grating edges reflecting 10 of the peak bandwidth ( 1 nm). The filter comprises three distinct sections, as shown in Fig. 5(a). In the central section of the device, corresponding to the apodized CDC, the mode travelling to the right through the Huygens’ waveguide is coupled backwards to the SWG waveguide. The backward coupling occurs within a 10-nm wide band centered at = 1550 nm. The SWG waveguide segments are modulated based on the data from Fig. 4(d) and following the target Gaussian bandgap profile. The results of these calculations are shown in Fig. 5(b). After coupling to the SWG waveguide, the mode travels to the left and transitions into a 450-nm wide strip waveguide over a length of = 10 m. The strip waveguide then bends away from the Huygens’ waveguide over a length of = 35 m and using a bending radius of R=20 m. Identical structures are also included at the output of the SWG waveguide to extract any directionally coupled light. Experimental results of the filter are shown in Fig. 5(c). Bandwidth remained fixed at 10 nm after fabrication, and the power loss was around 1 dB when referenced to the straight Huygens’ waveguide transmittance. We attribute the observed blue-shift of the center wavelength to deviations in the layout feature sizes during fabrication. An interference pattern was observed at the add port, which results from residual directional coupling between the Huygens’ and the SWG waveguides. However, this interference had minimal impact on the transmittance at the through port.
To conclude this section, we experimentally assessed the behaviour of the designed apodized CDC when integrated with a racetrack resonator. The device under test is shown in the SEM images of Fig. 6(a). The coupling length was varied from 20 m to 300 m in several steps, which in turn modified the free spectral range (FSR) of the resonances from 2.3 to 0.5 nm, respectively. A more detailed view of the racetrack resonances is provided in Fig. S5 in Supplement 1. The transmission spectra of the CDC-loaded racetrack resonators at the through, add and drop ports are presented in Fig. 6(b-g). The observed resonances exhibited an exceptionally low FSR, while the CDC also limited the light coupling spectrum of the racetrack to a 10-nm rejection band centered within the Huygens’ operation bandwidth. This demonstrated the ability to tailor the Huygens’ racetrack resonances with using an envelope function provided by an apodized SWG-assisted CDC.
3 Conclusion
In summary, this work has introduced and experimentally validated a new class of integrated photonic components based on Huygens’-resonant microring resonators, and directional and contra-directional couplers, all engineered for operation at the 1550 nm telecommunication wavelength. By exploiting the unique dispersion and group velocity properties of Huygens’ waveguides, we demonstrated efficient evanescent coupling in micro-ring resonators implemented in both directional and contra-directional configurations, relevant for compact, high-performance add-drop filters. We also explored the broader implications of these structures for controlling group index and group velocity dispersion, pointing to promising applications in nonlinear and quantum photonics. A key innovation is the development of a hybrid subwavelength grating–Huygens’ contra-directional coupler, enabling backward coupling across resonant and non-resonant metawaveguides and supporting broad spectral rejection bandwidths. This advancement facilitates the design of FSR-free racetrack resonators, opening new pathways for spectral engineering. Altogether, we anticipate that this work will open new research avenues, enabling incorporation of Huygens’ waveguide in the future integrated photonic platforms for both classical and quantum optical communication, computing and sensing applications.
Funding The project was funded by the (1) National Research Council Canada (NRC)’s (i) Postdoctoral Fellowship Program, (ii) Collaborative Science, Technology and Innovation Program (CSTIP) Small Teams (ST-R2-01-02), and by the (2) Deutsche Forschungsgemeinschaft (DFG, German Research Foundation, project number 448835038).
Acknowledgment M. Saad Bin-Alam acknowledges the support from the NRC’s Postdoctoral Fellowship in Metamaterial Integrated Photonics, and is thankful to Dr. Shurui Wang and Martin Vachon for their help in the experimental characterizations.
Disclosures The authors declare no conflicts of interest.
Data Availability Statement Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.
Supplemental document See Supplement 1 for supporting content.